# IDs for AI Systems

### v1

@@ 4 Demand for ID Use
IDs seem useful for actors engaging with an AI system, but why would a user or deployer of that AI system choose to present an ID? In short, governments, service providers, and parties interacting with AI systems all have interests in ID use, as well as means to incentivize (or mandate) it.
@@ 4.1 Governments
Interests in ID use: Governments may wish to disincentivize and investigate harms caused through AI systems. In the case of a financial scam for instance, the IDs of AI systems involved in any transactions could aid investigation. Similarly, governments have an interest in preventing fraud, and therefore already mandate that certain financial transactions involve ID checks.
Mandating ID use: Governments could mandate that certain, highly consequential service providers obtain IDs from any AI systems with which they interact. For example, financial institutions could request IDs when AI systems make large financial transactions.
@@ 4.2 Service Providers
Interests in ID use: IDs could be advantageous for service providers. First, IDs could help to reduce service abuse. Plugins and services—as well as the unique vulnerabilities of AI systems, such as prompt injection attacks (Zhan et al. 2024; Greshake et al. 2023)—could allow attackers to harm users or the service provider (Stumpp 2023). As a countermeasure, service providers could use IDs to block abusive instances. Since attackers could avoid block lists by creating instances with new IDs, service providers could limit the privileges of instances they have not interacted with before, and provide preferential services to certain, trusted IDs (e.g., instances with a history of good behavior; see more below). IDs could also aid own incident investigation processes. Second, service providers could use IDs to adapt services to the usage patterns of particular instances, which could help to retain service users and reduce costs.
Incentivizing ID use: Service providers could develop, or encourage the development of, plugins that require IDs. AI systems can already perform a variety of useful tasks through plugins, including web searches, email communication, and stock trading (OpenAI 2023a; Richards 2023; Wu et al. 2023; Anthropic 2024b). In comparison to an AI system mimicking a human’s interaction with a computer (e.g., interacting with individual web page elements), plugins may be more reliable and performant for users. AI systems continue to have difficulty with the former (Furuta et al. 2024; Tao et al. 2023; Gur et al. 2024, 2023; Xie et al. 2024), but plugins already enable useful tasks. As well, a lack of constraints with such mimicry—e.g., accidental file deletion from mouse/keyboard operation—could result in more severe problems than with plugins, which constrain the actions available to the AI system.
Yet, any actor could write a plugin for a service that does not require IDs. A potential response could be for service providers to restrict their services in the absence of an ID. Restrictions—such as rate limits, throttling, or quotas—already reduce abuse of many existing services (Cloudflare 2024). The design of restrictions could still allow humans to interact with services while limiting AI systems without IDs to human speeds, so as to reduce the impact of potential abuse. On the other hand, service providers could provide additional privileges to AI systems with IDs. This selectivity could incentivize ID use, since systems that presented IDs would have performance advantages.
Other actors could bypass plugins in any case by having an AI system mimic human interaction with a computer. A potential countermeasure could be CAPTCHAs-like methods to distinguish between human and non-human actions. Although software systems have become increasingly capable of bypassing CAPTCHAs, [^7] Task-based CAPTCHAs are no longer as common as CAPTCHAs based on behavior patterns, such as how a mouse is moved across the screen (Shet 2014). Yet, AI systems may one day be able to imitate human behaviour patterns as well (Adept 2022). they could be an interim option in anticipation of more robust proof-of-personhood protocols (Buterin 2023). Entities that failed CAPTCHAs could be subject to ID requirements, or else face service restrictions.
@@ 4.3 Other Parties that Interact with AI Systems
Interests in ID use: Some parties, such as businesses, may only wish to interact with certain, trusted AI systems when performing transactions. IDs could provide the information to identify such systems. On the other hand, it is unclear if individual consumers would use IDs. Yet, future software—including users’ personal AI assistants (Gabriel et al. 2024)—could inspect IDs and reject potentially unsafe interactions (e.g., with jailbroken systems), similar to how browsers warn users before visiting websites without valid HTTPS certificates.
Incentivizing ID use: Large corporations could exert strong pressure for ID usage, through reducing or avoiding engagement with AI systems without IDs. Future software that inspects IDs could also exert similar pressure. Analogously, websites that do not use HTTPS are disadvantaged since popular web browsers warn users when they are making an insecure connection.
@@ 4.4 Where to Require IDs
While IDs could be useful as described above, universal ID requirements might not be warranted. As we will discuss in Section 7, ID implementation imposes burdens and risks, and there remains uncertainty about the broader impacts of ID usage. As such, if governments or service providers decide to incentivize or require ID use, they should target where IDs could be most useful, such as high-stakes settings. We provide some examples of high-stakes tasks which could justify IDs.
Making large financial transactions: Some service providers could allow AI systems to perform financial transactions, such as making purchases, trading stocks (khan 2024), or otherwise transferring funds. Such service providers could require IDs for transactions involving significant sums. IDs could help to reduce the risk of a variety of negative outcomes, by enabling service providers to filter out potentially untrustworthy AI systems (e.g., those that may likely malfunction and make mistaken transactions) and investigate potential incidents.
Healthcare interactions: For better or for worse, some users are already turning to chatbots for health advice (Lucas 2024). IDs could help users make informed choices about such interactions. If AI systems mediate increasingly many medical interactions (Singhal et al. 2023), IDs could become more useful for establishing trust. In response to current trends, a variety of organizations—such as medical device manufacturers and hospitals—could provide plugins to connect AI systems to users’ healthcare data. Patients could have personal assistants that communicate with the AI systems of healthcare providers, so as to facilitate diagnosis and treatment.
5 ID Implementation for Centralized Deployments
We assess a basic ID system that deployers of AI systems could implement. Our discussion in this section will be agnostic to the applications where IDs are used (see Section 4.4 for discussion of where ID use may be warranted). In Section 6, we explore how IDs could work in more decentralized deployments.
@@ 5.1 Assigning IDs
Reloaded instances: Suppose that a deployer saves instances’ states (i.e., the history of interaction) and provides separate functionality for users to reload them. We tentatively suggest that deployers retain the ID of the original instance for the reloaded instance, since some attributes (e.g., prior incidents, context) of the original instance could be relevant for understanding the behaviour of the reloaded instance. Alternatively, if a deployer does decide to assign a new identifier to the reloaded instance, a link to the ID of the original instance would maintain accessibility to the relevant attributes.
@@
Composite systems: Software frameworks could give users the ability to create new instances (a composite) out of a collection of existing instances (Wu et al. 2023). If deployers run such software frameworks for users, they could detect the formation of composites. Since the constituent instances can continue running alongside the composite, composites should have new IDs, but could have their IDs linked to their constituents’.
Fine-tuned systems: Some deployers provide fine-tuning functionality to users. Since fine-tuning can change the behaviour of a system, a fine-tuned system $A^{\prime}$ should have different system_identifier than the original system $A$, and link to $A$’s system_identifier . Although it may be useful to include some documentation about the fine-tuning data in an ID, user privacy expectations could complicate such inclusion. Some deployers monitor fine-tuning data only for abuse detection and legal compliance (ChrisHMSFT et al. 2023; OpenAI 2024a); inclusion of fine-tuning data in an ID likely requires stronger justification.
@@ 5.3 Accessibility
To help to ensure that secondary parties maintain access to the ID, the deployer could include IDs in watermarks and metadata of media outputs (i.e., text, images, video). Yet, it is unclear how robust such methods may be to user removal (Zhang et al. 2023; OpenAI 2024c).
@@
5.4 Verifiability
To ensure verifiability, we explore an approach similar to that of digital certificates:
  1. (1)
    When an instance is created (e.g., when a new chat session starts), the deployer generates an ID for the instance.
  2. (2)
    Whenever the instance generates an output, the deployer attaches a digital signature for the combined ID and the output (e.g., via hashing the output).
  3. (3)
    The deployer sends the ID, signature, and output to the party interacting with the instance.
We emphasize that an ID would be associated to outputs via digital signatures. Hence, each output would have a different signature. Even though the information contained an ID might not change between different outputs, the signature verifying that the ID corresponds to an instance’s output would. This step is crucial to prevent instance spoofing. Otherwise, any actor could copy an ID, generate an output from an arbitrary instance, and present the copied ID as that of the arbitrary instance. [^9] Note that the signatures on HTTPS certificates do not need to be based on website content. Even if an attacker copied the details of an HTTPS certificate, they would not have the private key necessary for establishing a secure connection with a user. A digital signature also makes tampering and ID spoofing computationally difficult.
Deployers could sign outputs at differing granularities. For example, given a text output, should a deployer generate a signature for each token of the output? As a general principle, we suggest that the to-be-signed output should at least contain a functional unit for the party receiving the ID. For example, suppose a service provider receives requests in a JSON string. A signature of the ID and closing brace (“}”) of the JSON string would allow the attacker to replace the payload of the JSON string while maintaining the validity of the signature. On the other hand, a signature of an ID and the entire JSON payload would avoid this failure mode. We explore some additional examples of applying this principle:
  • Users: Deployers could sign the outputs generated since the last user input.
  • Other parties interacting with the instance: The deployer could sign the output which it expects the party to see. For example, if deployers provide a way for different instances to communicate with each other, the deployer should sign the sent messages.
However, requiring signature of the output comes with a crucial caveat. If any actor were to change or further process the output of an instance, the digital signature would not be valid for the modified output. Much flexibility could be lost under this implementation, if IDs were to be required in many settings.
Part of preventing ID spoofing involves guaranteeing that the digital signature indeed belongs to the deployer. Existing public key infrastructure for doing so involves a trusted third party known as a certificate authority (CA), who certifies that public keys (which allow third parties to verify digital signatures) belong to stated parties (the deployers in this case). The existing system of CAs could support the authenticity of the deployer’s signature. Alternatively, a collection of civil society and industry organizations could act as an initial CA, and delegate authority to certain deployers after having verified their identities. Future work should explore these options further, as well as decentralized alternatives to CAs that avoid concentrated points of failure (Sporny et al. 2022, 2024).
Figure 2: This figure illustrates the process in Section 5.4 of creating IDs that are verifiable.
6 ID Implementation for Decentralized Deployments
In some settings where ID use may be warranted (e.g., high-stakes settings; see Section 4.4), users themselves may run AI systems. We assess an implementation of IDs without deployers as an intermediary, which would require more involvement from users, CAs, and supporting software. As in Section 5, we adopt system_identifier:instance_identifier as the identifier.
6.1 A Basic Design
This design roughly follows that in Section 5, except that we will have AI frameworks—software tools such as the transformers library (Wolf et al. 2020), AutoGen (Wu et al. 2023), and AutoGPT (Richards 2023) that facilitate running AI systems—facilitate ID implementation. Similarly, automatic certificate issuance from Let’s Encrypt played an important role in the widespread adoption of HTTPS (Encrypt 2024). AI frameworks will require some additional functionality that does not yet exist, as we explain below.
  1. (1)
    The user registers for an account with a CA.
  2. (2)
    The AI framework lets a user log into their CA account.
  3. (3)
    Whenever a user starts an instance with an AI framework, the framework requests an identifier from the CA.
  4. (4)
    For each unit of output from the instance, the AI framework encrypts the output and sends it to the CA.
  5. (5)
    The CA produces a digital signature for the combined identifier and hashed output.
  6. (6)
    The AI framework makes the ID and signature accessible with the output.
As in Section 5.4, signature of the output would prevent attackers from attaching a given ID to arbitrary instances. To ensure user privacy, the CA would only sign a hashed output. Considerations for which outputs to sign (e.g., signing the entire JSON string of a service request) from Section 5.4 apply here as well.
As a potential way to ensure that CAs do not sign arbitrary outputs from users (e.g., those not even generated by an AI system), CAs could stipulate the use of approved AI frameworks. Techniques such as remote attestation and code signing could be useful for such requirements. [^10] Such techniques would ideally only apply to the loading and use of relevant library packages, rather than all parts of a user’s code base. Yet, only permitting the use of approved frameworks could have negative consequences, especially if IDs were required in many settings. Users would not be able to modify frameworks as they see fit. Bottlenecks in the approval process could also mean that users are pressured to use outdated frameworks, or frameworks that do not provide desired functionality.
Figure 3: This figure illustrates a potential process for ensuring verifiability of identifiers in decentralized deployments. See further details in Section 6.1.
Information to include: Some AI frameworks, such as HuggingFace, integrate with repositories or marketplaces where users can download different AI systems for use. Frameworks could link to the documentation that is often included in a system’s repository. However, unlike in Section 5.2, there is no guarantee that a repository would contain a system card or an incident database. If the AI framework provides a dedicated function for an ancestor instance to create descendant instances, the AI framework could establish ID links between the ancestor and the descendant. Otherwise, such links seem difficult to establish.
Accessibility: An AI framework has a similar role to that of a deployer in Section 5.3. The framework should ensure that any outputs of the system, including audiovisual outputs and queries to services, should contain the ID, whether embedded as a watermark or as metadata.
Verifiability: Verifiability would apply to the identifier, but not necessarily to other information attached to the ID. On HuggingFace for example, any anonymous user can upload a system with arbitrary documentation. Even so, such lack of verifiability may encourage parties to interact with instances of well-known systems, rather than instances of systems uploaded by anonymous users. This preference could incentivize users to use well-known systems and weaken the need for additional verifiability mechanisms.
6.2 Assigning IDs
In theory, AI frameworks could detect instance creation similarly to how deployers could in Section 5.1, and request new identifiers from CAs as appropriate. The main practical limitation is the flexibility that users have—by virtue of writing and running their own code—to create instances in arbitrary ways. It may not be feasible for an AI framework’s developer to account for all potential instance creation methods in advance, and decide how to assign IDs accordingly.
@@ 7.1 Failure to Inform
Parties could ignore IDs, just as internet users sometimes ignore indicators of untrustworthiness (Norris and Brookes 2021). Software—or future AI assistants (Gabriel et al. 2024)—could make informed decisions for users in some circumstances, just as browsers reject websites with invalid HTTPS certificates. However, it remains unclear what information would justify rejecting an interaction with an AI system. Other parties could misunderstand or misinterpret an ID’s attributes. For example, it can difficult to assess the external validity of evaluation results (Fourrier et al. 2023; Weidinger et al. 2023).
Moreover, ID attributes could be inaccurate. For example, deployers could unintentionally link to an incorrect version of system documentation, or may simply fail to update the documentation when a system is updated. OpenAI has occasionally released new versions of GPT-4, [^11] See https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4. but has not updated the GPT-4 system card (OpenAI 2023b) since the initial release, as of April 2024. Regular audits (Raji et al. 2020; Sharkey et al. 2023) could verify the accuracy of attributes.
@@ 7.2 Bypassing IDs
7.3 Limited Adoption
In decentralized deployments, the absence of supporting AI frameworks could hinder users from implementing IDs. Without such software, providers may be hesitant about requiring IDs if they wish to encourage use of their services. Consequently, there may not be enough demand to spur the development of supporting software. One way to resolve these problems is to focus on soft ID incentives with well-resourced actors. For example, service providers could experiment with ID incentives for companies, who could pilot initial implementations of supporting AI frameworks.
In the implementations we discussed, verifiability requires a signature on both the ID and the output. Modification of the output would hence invalidate the ID. Users could be hesitant to adopt IDs given this lack of flexibility. As well, many generated IDs could become invalidated simply because users unintentionally modify the outputs (e.g., a user accidentally removes some punctuation from generated text). ID requirements could be less onerous when outputs are unlikely to be modified, such as requests from an AI system to a service. Yet, absent technical implementations that can ensure verifiability while permitting modification, this limitation suggests that any initial ID requirements should be minimal.
7.4 Privacy and Security Risks
@@
In the context of the technical implementations we discussed, the trustworthiness and security of the CAs limits the verifiability of IDs. Apart from external threats, high-stakes situations could create incentives for collusion between authorities and key owners. With a compromised CA, a malicious actor could issue IDs for any AI system. For example, an attacker could assume the ID of a trusted AI system so as to interact with certain parties or obtain higher privileges with a service.
Decentralized alternatives to digital signature management could help to avoid these problems. Specifically, smart contracts deployed on distributed ledgers could potentially issue digital signatures without the need for a centralized authority (Omote 2023). The smart contracts would be publicly verifiable and are immutable as long as the integrity and availability of the distributed ledger as a whole, in contrast to a single trusted authority, is ensured. Yet, widespread adoption of such approaches is currently limited. Challenges include a lack of usable and tested software products, scalability and compute demands, and technical issues inherent in real-world implementations of distributed ledgers.
@@ 7.5 Broader Societal Consequences
Since IDs could enable anybody to identify and make decisions about particular instances, their use could have broad societal consequences.
IDs with additional user information could enable potentially harmful ranking of users. Actors implementing IDs may wish to include such information if it is correlated with undesirable outcomes, such as fraud. Yet, the existence of confounders could lead to instances of some users being unfairly denied access to services, similar to how toxicity classifiers can biased against African American English (Sap et al. 2019). Even if the relationship between information and undesirable outcome was causal, taking such information into account could still be unjustly discriminatory.
IDs could also enable influence over particular instances and their users. Even if an ID does not contain information about an instance’s interactions, external actors could record particular IDs have been involved in interactions, just as cookies can track user activity across websites. Instances that are active for long periods of time could reveal much information about a user’s preferences. Actors interacting with instances could try to influence their behaviour through means such as prompting. Advertisers and businesses could try to get instances to purchase or recommend their products or services for their corresponding users. Political actors could try to get instances to influence the voting behaviour of their users. Governments could attempt to enforce rules upon particular instances. Such influence could be strong, especially if personal AI assistants become much more widespread and central to daily interactions (Gabriel et al. 2024). The specific impacts of this influence deserve further study.
Finally, IDs could create a separate digital channel for interactions between trusted AI systems. The vast majority of digital interactions could in the future be between AI systems, especially if future AI assistants handle most of a user’s activities (Gabriel et al. 2024). If IDs enable reliable interactions, users may prefer their AI systems to interact only with AI systems that present IDs. Much digital activity could take place between networks of ID-bearing AI instances that trust (and can verify) each other. Entities that trust each other may be able to engage in more productive interactions. As a corollary, instances without IDs may become severely disadvantaged, relative to instances within ID-bearing networks. To avoid such disadvantage, users may have no choice but to submit to the potential negative consequences of IDs we have discussed heretofore in this section. On balance, the consequences of a separate digital channel for AI systems remain unclear.

### v3

@@ 5 Demand for ID Use
IDs seem useful for actors engaging with an AI system, but why would a user or deployer of that AI system present an ID? In short, governments, service providers, and parties interacting with AI systems all have interests in ID use, as well as means to incentivize (or mandate) it.
@@ 5.1 Governments
Interests in ID use: Governments may wish to disincentivize and investigate harms caused through AI systems. In the case of a financial scam for instance, the IDs of AI systems involved in any transactions could aid investigation (see Section 4.3). Similarly, governments have an interest in preventing fraud, and therefore already mandate that certain financial transactions involve ID checks.
Mandating ID use: Governments could mandate that certain, highly consequential service providers obtain IDs from any AI systems with which they interact. For example, financial institutions could request IDs if AI systems make large financial transactions.
@@ 5.2 Service Providers
Interests in ID use: IDs could aid incident investigation processes (see Section 4.3) and ensure that only trusted agents interact with services (see Sections 4.1 and 4.2), thereby disincentivizing or reducing service abuse.
Incentivizing ID use: Service providers could develop, or encourage the development of, plugins that require IDs. AI systems can already perform a variety of useful tasks through plugins, including web searches, email communication, and stock trading (OpenAI, 2023a; Richards, 2023; Wu et al., 2023; Anthropic, 2024b). In comparison to direct interaction with a user’s computer (e.g., interacting with individual web page elements), plugins could be more reliable and performant. AI systems continue to have difficulty with the former (Furuta et al., 2024; Tao et al., 2023; Gur et al., 2024; 2023; Xie et al., 2024), but plugins already enable useful tasks. As well, plugins could constrain the AI system’s actions in a more targeted way, so as to improve safety.
Yet, any actor could write a plugin that does not require IDs. A potential response could be for providers to restrict the services themselves, rather than just the plugins, in the absence of an ID. Restrictions, such as rate limits, could still allow humans to interact with services while reducing the damage an AI system could do.
For AI systems that interact with their users’ computers directly, CAPTCHAs-like methods could help to incentivize ID use. Entities that failed CAPTCHAs could be subject to ID requirements, or else face service restrictions. Although software systems have become increasingly capable of bypassing CAPTCHAs, [^9] Task-based CAPTCHAs are no longer as common as CAPTCHAs based on behavior patterns, such as how a mouse is moved across the screen (Shet, 2014). Yet, AI systems may one day be able to imitate human behaviour patterns as well (Adept, 2022). they could be an interim option in anticipation of more robust proof-of-personhood protocols (Borge et al., 2017).
Figure 3: Various actors have potential (dotted lines) methods to incentivize the use of IDs, whether directly or through using other actors.
@@ 5.3 Other Parties that Interact with AI Systems
Interests in ID use: Some parties, such as individuals or businesses, may only wish to interact with certain, trusted AI systems (see Section 4.2). Future software—including users’ personal AI assistants (Gabriel et al., 2024)—could inspect IDs and reject potentially unsafe interactions (e.g., with jailbroken systems), similar to how browsers warn users before visiting websites without valid HTTPS certificates.
Incentivizing ID use: Large corporations could exert strong pressure for ID usage, through reducing or avoiding engagement with AI systems without IDs. Future software that inspects IDs could also exert similar pressure. Analogously, websites without HTTPS are disadvantaged since popular web browsers warn users about insecure connections.
@@ 5.4 Where to Require IDs
While IDs could be useful, universal ID requirements might not be warranted. As we will discuss in Section 7, ID implementation imposes burdens and risks, and there remains uncertainty about the broader impacts of ID usage. As such, if governments or service providers decide to incentivize or require ID use, they should target settings where IDs could be most useful, such as when AI systems use services that have a direct impact on the world. We provide some examples of such services.
Making large financial transactions: Some service providers could allow AI systems to perform financial transactions, such as making purchases, trading stocks (Khan et al., 2024), or otherwise transferring funds. Such service providers could require IDs for transactions involving significant sums. IDs could help to reduce the risk of a variety of negative outcomes, by enabling service providers to filter out potentially untrustworthy AI systems (e.g., those that may likely malfunction and make mistaken transactions) and investigate potential incidents.
Contacting humans: Potential services for contacting humans include crowdwork platforms, telephone services, and online messaging. The ability to contact humans and potentially convince them to take actions in the world could greatly magnify the impact of an AI system. As we discussed in Section 4.3, IDs could help to disincentivize and catch misuse of these services.
6 Potential ID Implementation
We assess a basic ID system that deployers of AI systems could implement. Our discussion in this section will be agnostic to the applications where IDs are used (see Section 5.4 for discussion of where ID use may be warranted). We also omit discussion of verifiability, given our presentation in Section 3.3 and Figure 2 of how existing technologies could enable verifiability.
@@ 6.1 Assigning IDs
Reloaded instances: Suppose that a deployer saves instances’ states (i.e., the history of interaction) and provides separate functionality for users to reload them. We tentatively suggest that deployers retain the ID of the original instance for the reloaded instance, since some attributes (e.g., prior incidents, context) of the original instance could be relevant for understanding the behaviour of the reloaded instance. Alternatively, if a deployer does decide to assign a new identifier to the reloaded instance, a link to the ID of the original instance would maintain access to the relevant attributes.
@@
Composite systems: Software frameworks could give users the ability to create new instances (a composite) out of a collection of existing instances (Wu et al., 2023). If deployers run such software frameworks for users, they could detect the formation of composites. Since the constituent instances can continue running alongside the composite, composites should have new IDs, but could have their IDs linked to their constituents’.
Fine-tuned systems: Some deployers provide fine-tuning functionality to users. Since fine-tuning can change the behaviour of a system, a fine-tuned system $A^{\prime}$ should have different system_identifier than the original system $A$, and link to $A$’s system_identifier . Although it may be useful to include some documentation about the fine-tuning data in an ID, user privacy expectations could complicate such inclusion. Some deployers monitor fine-tuning data only for abuse detection and legal compliance (ChrisHMSFT et al., 2023; OpenAI, 2024a); inclusion of fine-tuning data in an ID likely requires stronger justification.
@@ 6.3 Access
To help to ensure that secondary parties maintain access to the ID, the deployer could include IDs in watermarks and metadata of media outputs (i.e., text, images, video). Yet, it is unclear how robust such methods may be to user removal (Zhang et al., 2023; OpenAI, 2024c).
@@ 7 Limitations
7.1 Misuse of IDs
If IDs were prevalent but verifiability was not assured, misuse of IDs could be a significant problem. For example, an attacker’s agent could masquerade as somebody else’s agent to perform illicit actions, such as sending spam or carrying out cyberattacks. Overdependence on such IDs could lead to false accusations against innocent users, who may not have the resources or expertise to defend themselves and point out flaws in the ID system. Ideally, the compromise of an ID should be made clear to all that can access it. IDs should also likely not be the sole or conclusive source of evidence for incident investigations. Concerns about the failure of ID verifiability mirror similar concerns about the reliability of watermarks (Gleichauf & Geer, 2024).
7.2 Decentralized Operation of Agents
We have not discussed how IDs could be implemented for AI systems which users run themselves, without a deployer. If service providers decide to require IDs, users will need straightforward ID implementation methods. Otherwise, they could be locked out of services or be forced to use a deployer. A potential inspiration is Let’s Encrypt (Encrypt, 2024), which automated the process of HTTPS certificate creation for websites. A similar service could work with software for running AI systems (e.g., transformers (Wolf et al., 2020), AutoGen (Wu et al., 2023)) to allow users to request and attach IDs. We leave practical details for future work.
@@ 7.3 Failure to Inform
Parties could ignore IDs, just as internet users sometimes ignore indicators of untrustworthiness (Norris & Brookes, 2021). Software—or future AI assistants (Gabriel et al., 2024)—could make informed decisions for users in some circumstances, just as browsers reject websites with invalid HTTPS certificates. However, it remains unclear what information would justify rejecting an interaction with an AI system. Other parties could misunderstand or misinterpret an ID’s attributes. For example, it can difficult to assess the external validity of evaluation results (Fourrier et al., 2023; Weidinger et al., 2023).
ID attributes could also be inaccurate. For example, deployers could unintentionally link to an incorrect version of system documentation, or simply fail to update it. OpenAI has occasionally released new versions of GPT-4, [^11] See https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4. but has not updated the GPT-4 system card (OpenAI, 2023b) since the initial release, as of April 2024. As a potential mitigation, regular audits (Raji et al., 2020; Sharkey et al., 2023) could verify the accuracy of attributes.
@@ 7.4 Bypassing IDs
7.5 Privacy Risks
@@ 7.6 Broader Societal Consequences
IDs with additional user information could enable potentially harmful ranking of users. Actors implementing IDs may wish to include such information if it is correlated with undesirable outcomes, such as fraud. Yet, the existence of confounders could lead to instances of some users being unfairly denied access to services, similar to how toxicity classifiers can biased against African American English (Sap et al., 2019). Even if the relationship between information and undesirable outcome was causal, taking such information into account could still be unjustly discriminatory.
IDs could also enable influence over particular instances and their users. Even if an ID does not contain information about an instance’s interactions, external actors could record particular IDs have been involved in interactions, just as cookies can track user activity across websites. Instances that are active for long periods of time could reveal much information about a user’s preferences. Actors interacting with instances could try to influence their behaviour through means such as prompting. Advertisers and businesses could try to get instances to purchase or recommend their products or services for their corresponding users. Political actors could try to get instances to influence the voting behaviour of their users. Governments could attempt to enforce rules upon particular instances. Such influence could be strong, especially if personal AI assistants become much more widespread and central to daily interactions (Gabriel et al., 2024). The specific impacts of this influence deserve further study.
Finally, IDs could create a separate digital channel for interactions between trusted AI systems. The vast majority of digital interactions could in the future be between AI systems, especially if future AI assistants handle most of a user’s activities (Gabriel et al., 2024). If IDs enable reliable interactions, users may prefer their AI systems to interact only with AI systems that present IDs. Much digital activity could take place between networks of ID-bearing AI instances that trust (and can verify) each other. Entities that trust each other may be able to engage in more productive interactions. As a corollary, instances without IDs may become severely disadvantaged, relative to instances within ID-bearing networks. To avoid such disadvantage, users may have no choice but to submit to the potential negative consequences of IDs we have discussed heretofore in this section. On balance, the consequences of a separate digital channel for AI systems remain unclear.